OpenStudy (anonymous):

Simplify x^4-3x^3/x^2-9

OpenStudy (austinl):

Ok, I think I got this. \[\frac{x^4-3x^2}{x^2-9}\] For this particular problem, you would want to factor both the top and bottom. Do you think that you could do that for me? If You need help, I can certainly give it.

OpenStudy (anonymous):

yea i got this x-3/x+3 i just wasnt sure if i was right

OpenStudy (austinl):

That is not the answer that I got. For the top part of the function, we have \[x^4 -3x^3\] We can then take out an x^3 to get \[x^3(x-3)\] On the bottom we get (x+3)(x-3) so we now have \[\frac{x^3(x-3)}{(x-3)(x+3)}\] Then after we cancel out like terms, we have \[\frac{x^3}{(x+3)}\] Does that make sense?

OpenStudy (anonymous):

oooo ok yes i get it now :) thnxs

OpenStudy (austinl):

no problem